OpenStudy (ray1998):

Help please? Will give both fan and medal. Write and graph a direct variation equation that passes through the given point. (5, -6) Thank you in advance!

Nnesha (nnesha):

equation for direct variation is \[\rm y=kx\] where k is the constant of variation substitute x and y for the values solve for k (x,y)-->(5,-6)

OpenStudy (yumyum247):

just about what i was thinking!!! LOL good gob neesha!!

OpenStudy (ray1998):

Okay, thank you. How would I start to solve for K?

Nnesha (nnesha):

show me what did you get after replacing x and y with (5,-6) x=? y=?

OpenStudy (ray1998):

5 = k(-6) like that?

Nnesha (nnesha):

remember (x,y) order pair where first number is x and 2nd one y (5,-6) x= ? y=? remember the alphabetical order

OpenStudy (ray1998):

5 = x and -6 = y

Nnesha (nnesha):

yep right so substitute x for 5 and y for -6

OpenStudy (ray1998):

y = kx -6 = k(5)

Nnesha (nnesha):

looks good to solve for k you should move the number 5 to the opposite side how would you do that ? how would you cancel out 5 from right side ?

OpenStudy (ray1998):

By multiplying it by itself? That way, the 5 on the right side gets crossed out and then you bring the down while multiplying -6 by 5; which would give you your answer? -6 = k(5) 5 5

Nnesha (nnesha):

when you multiply 5 by itself what would you get ?

OpenStudy (ray1998):

0? Cause they cross out, which makes you drop the k. Right?

OpenStudy (yumyum247):